5) Find the area under the normal distribution curve, show your work:

a. to the left of z = -0.37

b. between z = 0.24 and z=-1.12

c. to the right of z = -1.92
Answer #1

solution: to the left of z = -0.37

p(z<-0.37)=0.3557    Using z table   

b. between z = 0.24 and z=-1.12

= P(-1.12 < Z <0.24)

= P(Z < 0.24) - P(Z <-1.12 )

Using z table   

=0.5948- 0.1314

=0.4634

c.

to the right of z = -1.92

p(z>-1.92)=1-p(z<-1.92)=1 -0.0274=0.9726     Using z table